Sol Summary – April 12th

Crew 279 Sol Summary Report 12-04-2023
Sol: 10
Summary Title: Sadly, close to the end…
Author’s name: Agnès Dekeyser, Crew Executive Officer
Mission Status: Nominal
Sol Activity Summary:
This morning the boys woke up early for their EVA. The Crew Scientist made some radioactivity measurements in the canyon with the help of the Commander while the GHO and Crew astronomers made some drone mapping of the area. They didn’t find any isotope but the mapping was a success.
In the meantime, the girls took some time to rest. They also cooked a delicious lunch with homemade hummus, pepper tapenade, and cheesy bread, for the great pleasure of the boys.
Right after lunch, a second team got prepared for their EVA to Kissing Camel Ridge W. They drove there and began to look for the cavities found by the GreenHab Officer on his 3D map made by drone mapping on EVA #4. They measured the cavities to check if the measurements were correct on the 3D map. They also found other cavities that didn’t appear on the 3D map. So, they enjoyed the time they had left to make another mapping. They also took the measurements of those cavities to make the same comparison as previously. Our Crew Executive Officer also took the opportunity to take some earth samples to see if she could find some martian microorganisms and compare their viability to one of her extremophiles.
The end of the afternoon was dedicated to everyone’s experiments before getting ready for the reports. Another productive day for our martionauts.
Look Ahead Plan:
Arriving at the end of the mission, our martionauts only have small things to take care of for their experiments. Tomorrow, the Crew Executive Officer will organise an EVA in order to collect her bacteria samples left in North Ridge and bring them back to the science dome where their viability will be assessed and compared to their unexposed analog. After that, the Crew Scientist will finish his radioactivity measurements around the MDRS. He still has 100 measurements to do around the MDRS and it should only take one morning to perform.
Apart from that the crew will continue their own experiments including the 3D printing of drone propellers and the analysis of their data.
Anomalies in work: nun-functional robotic observatory, the battery wires of Opportunity needs to be replaced, and issue with the battery of the suit 2.
Weather: partly cloudy, low 8°C/46°F, high 26°C/79°C
Crew Physical Status: Nominal
EVA: See above
Reports to be filed: Journalist report, Greenhab report, Operations report, EVA report, EVA request
Support Requested: None

Copyright © The Mars Society. All rights reserved. | Main Site